# Deblurring Neural Radiance Fields with Event-driven Bundle Adjustment

### v1

@@ 4.5.1. Quantitative Results of Reconstruction on Blender Scenes:
In Tables 3 and 4, we evaluate the reconstruction results of learned neural radiance fields in five blender scenarios from deblurring and novel views, respectively. SRN and SRN+NeRF achieve the best results in single-image-based deblurring methods. With the help of event data, D2Net is also slightly better than MPR. For Deblur-NeRF and BAD-NeRF, although they have been further improved compared to NeRF+D2Net, NeRF+MPR, and NeRF+SRN, they still cannot surpass event-enhanced E2NeRF and our proposed EBAD-NeRF. Note that there is no apparent difference between the E2NeRF and EBAD-NeRF on the Cozyroom scene where linear camera shake produces motion blur, and E2NeRF is even better than EBAD-NeRF on LPIPS. However, in the other four scenes with more complex camera shaking as described in Sec. 4.1, the results of EBAD-NeRF in both deblurring views and novel views are significantly better than E2NeRF, which indicates that lacking the optimization of poses limits the role of events in 3D reconstruction in E2NeRF, especially when facing complex camera motion. Accordingly, EBAD-NeRF uses event-driven bundle adjustment to jointly optimize the motion poses of the camera, further releasing the potential of event data in reconstructing a sharp NeRF with blurry images.
Table 6. Absolute Trajectory Error Results of Blender Scenes.
ATE
COLMAP-Blur (NeRF (Mildenhall et al., 2021)) 0.0476 $\pm$ 0.0171
COLMAP-EDI (E2NeRF (Qi et al., 2023)) 0.0419 $\pm$ 0.0159
Linear Bundle Adjustment (BAD-NeRF (Wang et al., 2023)) 0.0462 $\pm$ 0.0142
Event-driven Bundle Adjustment (Ours) 0.0383 $\pm$ 0.0143
4.5.2. Quantitative Results of Pose Estimation on Blender Scenes:
In Table 6, we use absolute trajectory error to evaluate the accuracy of camera motion trajectory estimation by different methods. The results in the table are the average of five blender scenes. COLMAP-Blur means directly inputting the blurry images into COLMAP to estimate the camera poses of each view. This is a typical pose estimation method in NeRF-related work. COLMAP-EDI represents the method used in E2NeRF. It uses the EDI (Pan et al., 2019) algorithm to first use events to deblur the image and then input the sharp image sequence into COLMAP to estimate camera poses of the blur process. Linear bundle adjustment represents the linear interpolation method used by BAD-NeRF to jointly learn the pose changes of each view. Our proposed event-driven bundle adjustment method introduces event into joint learning of camera pose changes and sharp NeRF, achieving the best results in restoring the camera pose, which also promotes EBAD-NeRF to further improve the reconstruction quantification results as in Tables 3 and 4.
4.5.3. Quantitative Reconstruction Results on Blender Scenes:
To verify the effectiveness of our method on real data, we conducted experiments with two real scenes captured by DAVIS-346. As shown in Table 5, the comparison results with other methods are highly consistent with the results on blender scenes in Tables 3 and 4, which verifies the reliability of our method in practical applications.
Figure 6. Qualitative results of deblurring views on blender Pool scene.
Figure 7. Qualitative results of novel views on blender Wine and Factory scenes (first and second row), real scene (third row).
@@ 4.6.1. Qualitative Results of Reconstruction on Blender Scenes
As shown in Figure 6 and the first and second row in Figure 7, we compared the results of the synthesis deblurring view on the pool scene and novel view on the Wine and Factory scenes, respectively. For the deblurring view results, the results of EBAD-NeRF are sharper than other state-of-the-art single-image deblurring methods, event-RGB-based deblurring methods, and deblurring NeRF methods. For the green plants and wooden boxes in the scene, some artifacts appear in the results of E2NeRF, and our method effectively recovers the texture details. In the Wine and Factory scenes, though E2NeRF realizes noticeable results for the characters and letters, our method is closer to ground truth in color, clarity, and line thickness.
Figure 8. Qualitative evaluation of trajectory fitting accuracy on Tanabata scene. The figure shows that the event-driven bundle adjustment significantly fits the camera motion closer to the ground truth in the roll, pitch, and yaw metrics than other methods. Note that though there is event enhancement on blurry image formation in E2NeRF, the motion trajectory is distorted and leads to the inaccurate NeRF.
@@ 4.6.2. Qualitative Results of Pose Estimation on Blender Scenes
In Figure 8, we compare the estimated camera trajectory of different methods from the three dimensions of roll, pitch, and yaw. The purple curve representing EBAD-NeRF is the closest to the dashed line representing ground truth, which is consistent with the quantitative data in Table 6.
4.6.3. Qualitative Results on Real Scenes
As in the third row of Figure 7, our method better recovers the white letters on the blackboard, compared to the pose-fixed event-RGB-based method E2NeRF, while BAD-NeRF is limited in reconstructing smooth areas (white areas on the podium). The results of other image-deblurring-NeRF methods and Deblur-NeRF are all limited by the severe blurry input, which is the same as the quantitative results in Table 5.
Experiments on synthetic and real data show that our method can learn better poses of the camera motion blur process with event-driven bundle adjustment compared to the previous deblurring NeRF method. On this basis, the event data is also superimposed on the image blur process supervision, eventually achieving better 3D reconstruction and novel view synthesis effects.
@@ 5. Conclusion
In conclusion, this paper presents a novel approach, leveraging event-driven bundle adjustment, to address the challenge of modeling camera motion blur within neural radiance fields (NeRFs). Our framework draws inspiration from the emerging field of event-based vision, which offers high temporal resolution and asynchronous events, ideal for capturing dynamic scenes with motion blur. By integrating intensity-change-metric event loss and photo-metric blur loss, our framework enables the simultaneous optimization of blur modeling alongside NeRF reconstruction. Experiments on synthetic and real-world datasets demonstrate the efficacy of our approach in accurately estimating camera poses and producing sharp NeRF reconstructions.

### v2

@@ 4.5. Quantitative Results
4.5.1. Quantitative Results of Pose Estimation on Blender Scenes:
We use the Evo tool (Grupp, 2017) to calculate the absolute trajectory error, considering both rotation and translation. The results in Table 3 are the average of five blender scenes, which shows the accuracy of camera motion trajectory estimation by different methods. COLMAP-Blur means directly estimating the camera poses with blurry images and COLMAP for each view. COLMAP-EDI represents the pose estimation method used in E2NeRF, which uses the EDI (Pan et al., 2019) algorithm to first deblur the image with events and then input the sharp image sequence into COLMAP to estimate camera poses during the blur process. BAD-NeRF learns the start and end poses of camera motion. It uses linear interpolation to simulate the intermediate poses and extends it with a cubic interpolation. The results of these two methods are also shown in the table. Our proposed event-driven bundle adjustment method introduces event into joint learning of camera pose changes and sharp NeRF, achieving the best results in restoring the camera pose, which also promotes EBAD-NeRF to further improve the reconstruction quantification results as in Sec. 4.5.2 and Sec. 4.5.3.
@@ 4.5.2. Quantitative Results of Reconstruction on Blender Scenes:
We evaluate the reconstruction results of five blender scenarios on both deblurring views (Table 4) and novel views (Table 5). SRN and SRN+NeRF achieve the best results among single-image-deblurring-based methods and even surpass the Deblur-NeRF. With the help of event data, D2Net and D2Net+NeRF are slightly better than MPR and MPR+NeRF. BAD-NeRF with cubic interpolation is further improved over BAD-NeRF with linear interpolation and SRN+NeRF, though it is still worse than event-enhanced methods E2NeRF and our proposed EBAD-NeRF on both novel view and deblurring view.
Note that there is no apparent difference between the E2NeRF and EBAD-NeRF on the Cozyroom scene where blur is only caused by linear camera motion, and E2NeRF is even better than EBAD-NeRF on LPIPS. However, in the other four scenes with more complex camera shaking as described in Sec. 4.1.1, the results of EBAD-NeRF in both deblurring views and novel views are significantly better than E2NeRF, which indicates that with fixed pre-estimated poses limits the effect of events in 3D implicit learning in E2NeRF, especially when facing complex camera motion. Accordingly, EBAD-NeRF uses event-driven bundle adjustment to jointly optimize the motion poses of the camera, further releasing the potential of event data in reconstructing a sharp NeRF with blurry images.
4.5.3. Quantitative Results of Reconstruction on Real Scenes:
To verify the effectiveness of our method on real data, we conducted experiments on real data mentioned in 4.1.2. As shown in Table 6, the comparison results with other methods are highly consistent with the conclusions obtained in Tables 4 and 5, which verifies the reliability of our method in practical applications.
@@ 4.6.1. Qualitative Results of Reconstruction on Blender Scenes
We evaluate the reconstruction results of the deblurring view on the Pool scene (Figure 6) and the novel view on the Wine and Factory scenes (first two rows in Figure 7). Since image-deblurring methods cannot synthesize novel view images, we do not show their results in Figure 7). The results of EBAD-NeRF are sharper than all other methods on both deblurring and novel views. For the green plants and wooden boxes in the Pool scene, some artifacts appear in the results of E2NeRF, and our method effectively recovers the texture details. In the Wine and Factory scenes, though E2NeRF realizes noticeable results for the characters and letters, our method is closer to ground truth in color, clarity, and line thickness.
Figure 6. Qualitative deblurring views rendering results of blender Pool scene.
Figure 7. Qualitative novel views rendering results of blender scenes (first and second row) and real scenes (third row).
Figure 8. Qualitative evaluation of trajectory estimation on Tanabata scene. EBAD-NeRF significantly fits the camera motion closer to the ground truth in the roll, pitch, and yaw metrics than all other methods. Although E2NeRF uses pre-deblurred images with events, the estimated motion trajectory is still distorted, leading to reconstruction degradation even with event enhancement during training.
4.6.2. Qualitative Results of Reconstruction on Real Scenes
As in the third row of Figure 7, our method better recovers the white letters on the blackboard compared to E2NeRF. BAD-NeRF is limited in reconstructing smooth areas (white areas on the podium). The results of other image-deblurring-NeRF methods and Deblur-NeRF are all limited by the severe blurry input, which is consist with the quantitative results in Table 6.
@@ 4.6.3. Qualitative Results of Pose Estimation on Blender Scenes
In Figure 8, we compare the estimated camera trajectory of different methods from the three dimensions of roll, pitch, and yaw. The purple curve representing EBAD-NeRF is the closest to the ground truth dashed line, which is consistent with the results in Table 3.
Experiments on synthetic and real data both show that our method can learn better poses of the camera motion blur process with event-driven bundle adjustment compared to the previous deblurring NeRF methods. Upon this basis, the event data is also superimposed on the image blur process supervision, eventually achieving better 3D reconstruction and novel view synthesis effects.
@@ 5. Conclusion
In conclusion, this paper presents a novel approach, leveraging event-driven bundle adjustment, to address the challenge of modeling camera motion blur within neural radiance fields (NeRF). We draw inspiration from the emerging field of event-based vision, which offers high temporal resolution events, ideal for capturing dynamic information in motion blur. By integrating intensity-change-metric event loss and photo-metric blur loss, our framework enables the simultaneous optimization of blur modeling alongside NeRF reconstruction. Experiments on synthetic and real-captured datasets demonstrate the efficacy of our approach in accurately estimating camera poses and producing sharp NeRF reconstruction results.
Appendix A Supplementary Qualitative Comparison on Synthetic Scenes
As shown in Figure 9, in the Cozyroom scene there is no apparent qualitative difference in the results of the BAD-NeRF (Wang et al., 2023), E2NeRF (Qi et al., 2023) and EBAD-NeRF because the camera motion is even and slight. However, in the other four scenes, the results of our EBAD-NeRF have been significantly improved compared to BAD-NeRF and E2NeRF, proving that when facing complicated camera motion, our proposed event-driven bundle adjustment can effectively provide accurate estimated poses for the NeRF network and with event-enhanced imaging simulation our method can recover a sharp NeRF from the severely blurry images and corresponding events.
In the Factory scene, EBAD-NeRF’s reconstruction of scars on the wall is closest to ground truth. In comparison, the results of E2NeRF are too thick, and the results of BAD-NeRF are still blurry. In the Pool scene, E2NeRF misses the lotus leaf in the pool and generates blurrier results in the Tanabata scene than EBAD-NeRF. Artifacts appear on the white flag in the results of E2NeRF in the Wine scene. Since there is no supervision of event data for pose estimation and image motion blur modeling, the results of BAD-NeRF are not sharp enough compared to ERGB-based methods E2NeRF and EBAD-NeRF. The pre-deblurring-based method MPR-NeRF (Zamir et al., 2021a), SRN-NeRF (Tao et al., 2018b), and D2Net-NeRF (Shang et al., 2021) is not good enough due to the limitations of the deblurring method itself. Overall, our EBAD-NeRF achieved the best results in all synthetic scenes.
Appendix B Supplementary Qualitative Comparison on Real-World Scenes
Figure 10 shows the two scenes of the real data. In the Bar scene, the synthesis novel view results of EBAD-NeRF are closest to ground truth, especially for the light spot in the upper left corner and the characters on the billboard. Other methods have obvious reconstruction errors of the light points and blur results of the characters. In the Classroom scene, our EBAD-NeRF obtains sharper letters and fewer artifacts in the synthesis results than other methods. The experiments on real-world scenes demonstrate that our EBAD-NeRF is more effective than the ERGB-based deblurring NeRF method E2NeRF, which without pose optimizing during the training. Also, our method has significantly improved over the image-based deblurring NeRF methods and NeRF with pre-deblurring images.
Appendix C Supplementary Video
We show the video rendering results for the ablation study, synthetic scenes, and real-world scenes on our project website: https://icvteam.github.io/EBAD-NeRF.html.
The video rendering results for the ablation study prove the effectiveness of our proposed event-driven bundle adjustment. For the synthetic scene, our method achieves the sharpest video rendering results. The results of E2NeRF are a little misaligned in space because the poses estimated by its pre-deblurring framework are slightly different from the initial poses given for the other methods. For the real-world scene, the results of NeRF and NeRF with pre-deblurring images are severely blurry and contain many cloudy materials. In comparison, our EBAD-NeRF has sharper rendering results with less cloudy materials than Deblur-NeRF, BAD-NeRF, and E2NeRF. Overall, our EBAD-NeRF achieves the best video rendering results in both real-world and synthetic scenes.
Figure 9. Supplementary Reconstruction Qualitative Comparison on Synthetic Scenes.
Figure 10. Supplementary Reconstruction Qualitative Comparison on Real Scenes.